How many grams of dry NH4Cl need to be added to 1.50 L of a 0.500 M solution of ammonia, NH3,to prepare a buffer solution that has a pH of 8.79? Kb for ammonia is 1.8*10^-5.
Answer #1
From Henderson Hesselbach equation :
       pOH = pKb + log [ conjugate base ] / [acid]                .............. (1)
Initial moles of  NH3 present in 1.50 L of 0.500 M = 1.50 L * 0.500 mol / L = 0.75 mol
     NH3 + H2O -----> NH4+ + OH-
     So,      p0H = pKb + log [ NH4+] / [NH3]
    14.00 - 8.79  = - log(1.8*10-5) + log ( x / 0.75 )         { where x is moles of NH4+}
                 5.21 = 4.75 + log ( x / 0.75 )
                 0.46 = log ( x / 0.75 )
             x / 0.75 = 100.46 = 2.88
            
Then, moles of NH4+ = 2.16 mol
Then mass of NH4Cl will be , 2.16 mol * ( 53. 49 g / 1mol ) = 116 g
Answer #2
From Henderson Hesselbach equation :
       pOH = pKb + log [ conjugate base ] / [acid]                .............. (1)
Initial moles of  NH3 present in 1.50 L of 0.500 M = 1.50 L * 0.500 mol / L = 0.75 mol
     NH3 + H2O -----> NH4+ + OH-
     So,      p0H = pKb + log [ NH4+] / [NH3]
    14.00 - 8.79  = - log(1.8*10-5) + log ( x / 0.75 )         { where x is moles of NH4+}
                 5.21 = 4.75 + log ( x / 0.75 )
                 0.46 = log ( x / 0.75 )
             x / 0.75 = 100.46 = 2.88
            
Then, moles of NH4+ = 2.16 mol
Then mass of NH4Cl will be , 2.16 mol * ( 53. 49 g / 1mol ) = 116 g
Answer #3
From Henderson Hesselbach equation :
       pH = pKb + log [ conjugate base ] / [acid]                .............. (1)
Initial moles of  NH3 present in 1.50 L of 0.500 M = 1.50 L * 0.500 mol / L = 0.75 mol
     NH3 + H2O -----> NH4+ + OH-
So, pH = pKb + log [ NH4+] / [NH3]
    8.79  = - log(1.8*10-5) + log ( x / 0.75 )         { where x is moles of NH4+}
    8.79 = 4.74 + log ( x / 0.75 )
          x = 8.4 * 103 mol
Then mass of NH4Cl will be , 8.4 * 103 mol * ( 53. 49 g / 1mol ) = 4.5*105 g